OpenStudy (anonymous):

What is common to the variation in the range and the height of a projectile? time of flight vertical acceleration horizontal acceleration vertical velocity

OpenStudy (shamim):

Range R=v0*cos theta not*v0*sin theta not*2/g R=v0^2*sin (2*theta not)/g

OpenStudy (shamim):

Height h=(v0*sin theta not)^2/(2g)

OpenStudy (shamim):

g is common in both R nd h

OpenStudy (shamim):

g=vertical acceleration

OpenStudy (shamim):

Response plz!!!!!!